Rx1nyDxyz:

Hii!! Could someone please help me with a math question? Thank you <33 Solve for x: −4x − 4 = −4(x + 2) 1. 0 2. −4 3. All real numbers 4. No solution I'm pretty sure it's 4, but I'd just like a second opinion :)

Vocaloid:

4 is incorrect. −4x − 4 = −4(x + 2) distributing the right side -4x - 4 = -4x - 8 If you subtract -4x from both sides, you get -4 = -8 which is clearly a false statement. Therefore, there is no x value that could possibly make this true, meaning there is no solution to this problem.